This is not an exact number in my book it is on a worksheet , “How long will it take to double my money?” At 8% interest rate and continuous compounding, what is the answer? Round your answer to the hundredth's place.? Please show how you got your answer
Answer by stanbon(48567)

At 8% interest rate and continuous compounding, what is the answer? Round your answer to the hundredth's place
A = Pe^(rt)
Doubling the money means A= 2P since P is the amount you start with.
2P = Pe^rt
2 =2^(0.08t)
Take the natural log of both sides to get:
ln2 = 0.08t
t= (ln2)/0.08
t=8.66 years(The money will double in 8.66 years)
Cheers,
Stan H.